# Selective “Selective Prediction”: Reducing Unnecessary Abstention in Vision-Language Reasoning

### v1

@@ Abstract
Prior work on selective prediction minimizes incorrect predictions from vision-language models (VLMs) by allowing them to abstain from answering when uncertain. However, when deploying a vision-language system with low tolerance for inaccurate predictions, selective prediction may be over-cautious and abstain too frequently, even on many correct predictions. We introduce ReCoVERR, an inference-time algorithm to reduce the over-abstention of a selective vision-language system without decreasing prediction accuracy. When the VLM makes a low-confidence prediction, instead of abstaining ReCoVERR tries to find relevant clues in the image that provide additional evidence for the prediction. ReCoVERR uses an LLM to pose related questions to the VLM, collects high-confidence evidences, and if enough evidence confirms the prediction the system makes a prediction instead of abstaining. ReCoVERR enables two VLMs, BLIP2 and InstructBLIP, to answer up to 20% more questions on the A-OKVQA task than vanilla selective prediction without decreasing system accuracy, thus improving overall system reliability.
@@
Tejas Srinivasan1,2   Jack Hessel3   Tanmay Gupta2   Bill Yuchen Lin2 Yejin Choi2,4   Jesse Thomason1   Khyathi Raghavi Chandu2 1University of Southern California  2Allen Institute for Artificial Intelligence 3Samaya AI  4University of Washington tejas.srinivasan@usc.edu, khyathic@allenai.org
@@ 1 Introduction
Instruction-tuned vision-language models (VLMs) Dai et al. (2023); Liu et al. (2023); Laurençon et al. (2023); Bai et al. (2023) have achieved strong accuracy on reasoning benchmarks, which typically require VLMs to produce an answer for each instance. For downstream use by people, however, these systems should abstain from answering when uncertain (e.g., by saying “I don’t know") Rajpurkar et al. (2018). Selective prediction systems De Stefano et al. (2000); El-Yaniv et al. (2010); Geifman and El-Yaniv (2017) aim to balance the number of predictions made (coverage) and the error rate on instances which do receive a prediction (risk).
However, a vanilla selective prediction system with low tolerance for incorrect predictions will abstain too frequently to be practical, even when the model answer may be correct Whitehead et al. (2022). For example, if a user specifies that the BLIP2 Li et al. (2023) predictions should be right at least 90% of the time, vanilla selective prediction will make a prediction for just 4% of A-OKVQA Schwenk et al. (2022) questions, with 94% of the correct predictions being abstained on.
@@
Our experiments on the A-OKVQA reasoning task using two VLMs (BLIP2 and InstructBLIP) demonstrate that ReCoVERR substantially increases the number of questions that a selective VLM system can answer, while keeping system risk under the specified risk tolerance, thus improving overall reliability (Section 5). ReCoVERR is particularly helpful for BLIP2, which has not been trained on the target task, by improving coverage by 20% and recall by 25-30%. Our analysis reveals that the ability to give accurate confidence estimates, and high estimates for correct predictions, is crucial to ReCoVERR’s performance. Further experiments demonstrate the importance of ensuring evidences are both reliable and relevant (Section 5.1), and that ReCoVERR tuned for a single task can be directly applied to new tasks without further tuning (Section 5.3). Our findings suggest that ReCoVERR is a promising solution towards building more reliable multimodal reasoning systems.
@@ 2 Multimodal Selective Prediction
We consider the task of answering a textual question about an image by drawing inferences from what is visually observed. A vision-language model (VLM) is given an input $x=(I,Q)\in\mathcal{X}$ consisting of an image $I$ and a question $Q$. It aims to produce an answer $a\in A$ from a closed or open set of possible answers.
@@
$\mathcal{S}_{\texttt{VLM}}(x)$ $=\begin{cases}a,&\text{if }g(a)=1\\ \varnothing,&\text{if not }g(a)=0\end{cases}$
Prior work (Whitehead et al., 2022) employs confidence-based selection, where instead of assuming access to a VLM which has been trained to abstain with explicit supervision, we assume a VLM that can produce both an answer, as well as a confidence score for that answer $\pi_{\texttt{{VLM}}}:\mathcal{A}\rightarrow[0,1]$ that estimates $\pi_{\texttt{{VLM}}}(a)=P(a|Q,I;\mathcal{M}_{\texttt{VLM}})$. Confidence-based selection relies on the confidence $\pi_{\texttt{{VLM}}}(a)$ being higher for correct answers than incorrect ones, on average. The decision function $g$ is a simple thresholding function, with a threshold $\gamma$: $g(a;\gamma)=\mathbbm{1}\{\pi_{\texttt{{VLM}}}(a)\geq\gamma\}$. We refer to this method as vanilla selective prediction.
@@ 2.1 Evaluating Selective Predictors
We evaluate selective prediction systems via coverage and risk El-Yaniv et al. (2010). Given a labeled evaluation dataset $\mathcal{D}=\{(I_{i},Q_{i},a_{i})\}_{i=1}^{N}$. Coverage is the the percentage of questions in $\mathcal{D}$ where the system chooses to make a prediction. Risk is the error rate on the questions where a prediction is made. For a $\gamma$-threshold selective prediction system, these metrics can be computed as:
@@ 3 ReCoVERR
NN: Maximum number of turns of evidence collection
@@ 13          $\mathcal{E}_{RR}\leftarrow\mathcal{E}_{RR}+[e_{j}]$
[…]
@@
Since ReCoVERR decides to make a prediction based on the collected evidences, the correctness likelihood of a given evidence is an upper bound for the correctness likelihood of the prediction. Since ReCoVERR aims to make predictions with correctness likelihood $\geq 1-r$, we only consider evidences $e_{j}$ that satisfy $\pi_{\texttt{{VLM}}}(a_{j})\geq 1-r$ as reliable for $\mathcal{E}_{R}$. ReCoVERR requires that the confidence estimate $\pi_{\texttt{{VLM}}}(a)$ is calibrated. A confidence estimate being calibrated means that for all predictions whose confidence is $\alpha\in[0,1]$, $\alpha$% of the predictions are actually correct. In our experiments, we first evaluate calibration of our VLMs’ confidence estimates (§ 4.2.1).
@@ 4 Experiments
4.1 Vision-Language Reasoning Task: A-OKVQA
@@ 4.2 Vision-Language Models
Figure 3: Calibration curves for BLIP-2 and InstructBLIP on A-OKVQA questions.
We experiment with two VLMs: BLIP2 Li et al. (2023) and InstructBLIP Dai et al. (2023), both using FlanT5-XL as the LLM backbone. [^3] https://github.com/salesforce/LAVIS InstructBLIP is instruction tuned on A-OKVQA; BLIP2 has not been trained on the A-OKVQA task.
@@ 4.2.1 Calibrated VLM Confidence Estimates
Confidence estimates for generative VLMs are not well defined. Straightforward estimates such as the product of answer token likelihoods can severely underestimate model confidence due to factors like surface form competition Holtzman et al. (2021). Inspired by Tian et al. (2023), we devise a Self-Prompting technique for estimating $\pi_{\texttt{{VLM}}}(a)$ by prompting the VLM to verify “yes” or “no” correctness of its predictions and examine the resulting probability distribution (full details in Appendix A). We further calibrated this confidence estimate using Platt scaling Platt (1999). We divide the A-OKVQA training data into 12,000 examples are used to fit the calibration model and 5,000 to evaluate calibration error. We find that off-the-shelf BLIP2 tends to be overconfident, evidenced by prediction confidence being consistently higher than accuracy (Figure 3). Platt scaling further reduces calibration error for both VLMs.
@@ 4.3 ReCoVERR Implementation Details
4.3.1 ReCoVERR Model-based Tools
In our experiments, the vision tools $\mathbb{M}_{\texttt{Vis}}$ consists of LVIS Gupta et al. (2019) for object detection, and Qwen-VL Bai et al. (2023) for region captioning. We use FlanT5-XL as a zero-shot sentence paraphraser $\mathcal{M}_{\texttt{QA}\rightarrow\texttt{S}}$ and entailment model $\mathcal{M}_{\texttt{NLI}}$. Since our VLMs use FlanT5-XL as the LLM backbone, our ReCoVERR instantiation does not require additional models. We use GPT-3.5 with temperature 1.0 for $\mathcal{M}_{\texttt{QGen}}$. See Appendix B for prompts.
4.3.2 ReCoVERR Hyperparameters
4.4 Selective Prediction Baselines
We compare ReCoVERR to a Vanilla Selective Prediction baseline Whitehead et al. (2022), where the model abstains if the VLM prediction confidence $g(a)=\{1\text{ if }\pi_{\texttt{{VLM}}}(a)\leq\gamma_{@r}\text{ else }0\}$, and to a Vision Tools baseline. For Vision Tools, image caption from the VLM, objects detected by LVIS, and region captions from Qwen-VL are provided as evidence to the NLI model $\mathcal{M}_{\texttt{NLI}}$ directly, with no rounds of evidence collection, equivalent to lines 1-5 of Algorithm 1 followed by lines 14-15.
4.5 Evaluation Metrics
We measure risk ($\mathcal{R}$) and coverage ($\mathcal{C}$) (Eqs 2, 1) across system predictions and abstentions. Additionally, we calculate Effective Reliability $\Phi_{c}$ Whitehead et al. (2022), a metric that rewards correct predictions, assigns a penalty $c$ to incorrect predictions, and assigns zero reward to abstentions. We assign a penalty of $c=1$ for incorrect answers ($\Phi_{1}$). Finally, we define Selective Prediction Recall ($\mathbb{R}_{\texttt{SP}}$), the percentage of correct answers ($\text{Acc}(a)=1$) in the dataset that were answered.
@@ 5 Results and Analysis
At 20% risk tolerance, we find that ReCoVERR improves coverage, effective reliability, and recall over both baselines while staying below or at the specified risk tolerance. At 10% risk tolerance, ReCoVERR and the Vision Tools baseline tend to slightly overshoot the risk tolerance (by 1-2% for InstructBLIP and 3-4% for BLIP2). We further see that ReCoVERR shows large improvements of up to 20% on coverage, up to 15% on effective reliability and up to 33% on recall for off-the-shelf BLIP2 and calibrated InstructBLIP.
Figure 4: Recall vs risk for different selective prediction methods applied to different VLMs, at two specified risk tolerances (10% and 20%).
Figure 4 visualizes how different selective prediction methods compare on risk $\mathcal{R}$ and recall $\mathbb{R}_{\texttt{SP}}$. We highlight several takeaways (T*). T1: For the same risk tolerance, vanilla selective prediction has lower coverage and recall for BLIP2 than InstructBLIP. BLIP2 has not been trained on A-OKVQA, unlike InstructBLIP, and therefore produces more uncertain answers. We further see that ReCoVERR results in larger improvements for BLIP2. T2: Off-the-shelf BLIP2 sees largest risk increase with ReCoVERR. The calibration curves in Figure 3 indicate that off-the-shelf BLIP2 makes overconfident estimates, leading to more unreliable evidences being collected and resulting in a large risk increase for ReCoVERR. T3: At 10% risk, most of the additional risk and recall is caused by Vision Tools, rather than ReCoVERR’s evidence collection. This finding is especially true for calibrated BLIP2 and off-the-shelf InstructBLIP. T4: At 20% risk tolerance, ReCoVERR’s evidence collection leads to greatest recall improvement for off-the-shelf BLIP2 and calibrated InstructBLIP.

### v2

@@ Abstract
Selective prediction minimizes incorrect predictions from vision-language models (VLMs) by allowing them to abstain from answering when uncertain. However, when deploying a vision-language system with low tolerance for inaccurate predictions, selective prediction may be over-cautious and abstain too frequently, even on many correct predictions. We introduce ReCoVERR, an inference-time algorithm to reduce the over-abstention of a selective vision-language system without increasing the error rate of the system’s predictions. When the VLM makes a low-confidence prediction, instead of abstaining ReCoVERR tries to find relevant clues in the image that provide additional evidence for the prediction. ReCoVERR uses an LLM to pose related questions to the VLM, collects high-confidence evidences, and if enough evidence confirms the prediction the system makes a prediction instead of abstaining. ReCoVERR enables three VLMs (BLIP2, InstructBLIP and LLaVA-1.5) to answer up to 20% more questions on the VQAv2 and A-OKVQA tasks without decreasing system accuracy, thus improving overall system reliability. Our code is available at https://github.com/tejas1995/ReCoVERR.
@@
Tejas Srinivasan1thanks:  Work done by Tejas as part of an internship at AI2.  Jack Hessel2  Tanmay Gupta3  Bill Yuchen Lin3 Yejin Choi3,4  Jesse Thomason1  Khyathi Raghavi Chandu3 1University of Southern California  2Samaya AI 3Allen Institute for Artificial Intelligence  4University of Washington tejas.srinivasan@usc.edu
@@ 1 Introduction
Instruction-tuned vision-and-language models (VLMs) Dai et al. (2023); Liu et al. (2023); Laurençon et al. (2023); Bai et al. (2023) have achieved strong accuracy on reasoning benchmarks, which typically require VLMs to produce an answer for each instance. For downstream use, however, these systems should abstain from answering when uncertain (e.g., by saying “I don’t know") Rajpurkar et al. (2018). Selective prediction systems De Stefano et al. (2000); El-Yaniv et al. (2010) aim to balance the number of predictions made (coverage) and the error rate on predicted instances (risk).
However, a vanilla selective prediction system with low tolerance for incorrect predictions will abstain too frequently to be practical, even when the model answer may be correct Whitehead et al. (2022a). For example, if a user specifies that the BLIP2 Li et al. (2023) predictions should be right at least 90% of the time, vanilla selective prediction will make a prediction for just 4% of A-OKVQA Schwenk et al. (2022) questions, with 94% of the correct predictions being abstained on.
@@
We experiment on the VQAv2 and A-OKVQA visual reasoning tasks using three VLMs: BLIP2, InstructBLIP and LLaVA-1.5. For all three VLMs, ReCoVERR substantially increases the number of questions answered by the selective VLM system, while keeping system risk under the specified risk tolerance (Section 5). ReCoVERR is particularly helpful for BLIP2, which has not been trained on the target task, by improving coverage by 20% and recall by 25-30%. Our analysis reveals that the ability to give accurate confidence estimates, and high estimates for correct predictions, is crucial to ReCoVERR’s performance. Further experiments demonstrate the importance of ensuring evidences are both reliable and relevant (Section 5.1), and that ReCoVERR tuned for a single task can be directly applied to new tasks without further tuning (Section 5.3). Our findings suggest that ReCoVERR is a promising solution towards building more reliable multimodal reasoning systems.
@@ 2 Multimodal Selective Prediction
We consider the task of answering a textual question about an image by drawing inferences from what is visually observed. A vision-language model (VLM) is given an input $x=(I,Q)\in\mathcal{X}$ consisting of an image $I$ and a question $Q$, and aims to produce an answer $a\in A$ from a closed or open set of possible answers.
@@
$\mathcal{S}_{\texttt{VLM}}(x)$ =$=\begin{cases}a,&\text{if }g(a)=1\\ \varnothing,&\text{if not }g(a)=0\end{cases}$
Prior work (Whitehead et al., 2022a) employs confidence-based selection, where instead of assuming access to a VLM which has been trained to abstain with explicit supervision, we assume a VLM that can produce both an answer, as well as a confidence score for that answer $\pi_{\texttt{{VLM}}}:\mathcal{A}\rightarrow[0,1]$ that estimates $\pi_{\texttt{{VLM}}}(a)=P(a|Q,I;\mathcal{M}_{\texttt{VLM}})$. Confidence-based selection relies on the confidence $\pi_{\texttt{{VLM}}}(a)$ being higher for correct answers than incorrect ones, on average. The decision function $g$ is a simple thresholding function, with a threshold $\gamma$: $g(a;\gamma)=\mathbbm{1}\{\pi_{\texttt{{VLM}}}(a)\geq\gamma\}$. We refer to this method as vanilla selective prediction.
@@ 2.1 Evaluating Selective Predictors
We evaluate selective prediction systems via coverage and risk El-Yaniv et al. (2010). Given a labeled evaluation dataset $\mathcal{D}=\{(I_{i},Q_{i},a_{i})\}_{i=1}^{N}$. Coverage ($\mathcal{C}$) is the percentage of questions in $\mathcal{D}$ where the system chooses to make a prediction. Risk ($\mathcal{R}$) is the error rate on the questions where a prediction is made. For a $\gamma$-threshold selective prediction system, these metrics are computed as:
@@ 3 ReCoVERR
$N$: Maximum number of turns of evidence collection
@@ 13                         $\mathcal{E}_{RR}\leftarrow\mathcal{E}_{RR}+[e_{j}]$
16      if $P(\mathcal{E}_{RR}\text{ entails }\mathcal{H};\mathcal{M}_{\texttt{NLI}})\geq\pi_{\texttt{NLIMin}}$ then
@@
Since ReCoVERR decides to make a prediction based on the collected evidences, the correctness likelihood of a given evidence is an upper bound for the correctness likelihood of the prediction. Since ReCoVERR aims to make predictions with correctness likelihood $\geq 1-r$, we only consider evidences $e_{j}$ that satisfy $\pi_{\texttt{{VLM}}}(a_{j})\geq 1-r$ as reliable for $\mathcal{E}_{R}$. ReCoVERR requires that the confidence estimate $\pi_{\texttt{{VLM}}}(a)$ is calibrated. A confidence estimate being calibrated means that for all predictions whose confidence is $\alpha\in[0,1]$, $\alpha$of the predictions are actually correct. In our experiments, we first evaluate calibration of our VLMs’ confidence estimates (§ 4.2.1).
@@ 4 Experiments
Figure 3: Calibration curves for BLIP-2, InstructBLIP and LLaVA-1.5 on A-OKVQA questions.
@@
4.1 Vision-Language Reasoning Tasks: A-OKVQA and VQAv2
A-OKVQA Schwenk et al. (2022) is a VQA task designed to require reasoning over external knowledge and commonsense alongside image-based information. VQAv2 Goyal et al. (2017) is a VQA task that requires reasoning about images, sometimes involving commonsense reasoning. We evaluate on the A-OKVQA validation set ($n=1,075$) and 1,000 examples from the VQAv2 validation set. Both tasks involve open-ended answer generation, without any choices provided to the model.
@@ 4.2 Vision-Language Models
We experiment with three VLMs: BLIP2 Li et al. (2023), InstructBLIP Dai et al. (2023), and LLaVA-1.5 Liu et al. (2023). Both BLIP models use FlanT5-XL as the LLM backbone. [^3] https://github.com/salesforce/LAVIS InstructBLIP and LLaVA-1.5 are instruction tuned on both VQAv2 and A-OKVQA; BLIP2 has not been trained on either task.
@@ 4.2.1 Calibrated VLM Confidence Estimates
One key requirement for ReCoVERR is that, for a prediction $a$, the VLM can also return a calibrated confidence score $\pi_{\texttt{{VLM}}}(a)\in[0,1]$. Thus, a first step when applying ReCoVERR to a new VLM is identifying a confidence function that produces calibrated confidence estimates.
Confidence estimates for generative VLMs are not well defined. Straightforward estimates such as the product of answer token likelihoods can severely underestimate model confidence due to factors like surface form competition Holtzman et al. (2021). Inspired by Tian et al. (2023), we devise a Self-Prompting technique for estimating $\pi_{\texttt{{VLM}}}(a)$ by prompting the VLM to verify “yes” or “no” correctness of its predictions and examine the resulting probability distribution (full details in Appendix A). The Self-Prompting confidence can be further calibrated using Platt scaling Platt (1999).
We compare three methods for extracting VLM confidence scores: product of token probabilities, mean of token probabilities, and Self-Prompting. For each VLM, we evaluate the calibration error of these confidence functions on a set of 5,000 examples from the A-OKVQA training data. The remaining 12,000 examples are used to calibrate Self-Prompting with Platt scaling. We find that Self-Prompting (both off-the-shelf and calibrated) has the lowest calibration error for the BLIP models, whereas the mean of token probabilities has lowest error for LLaVA-1.5 (Figure 3). We therefore use these respective methods for estimating VLM confidence in our experiments with ReCoVERR.
@@ 4.3 ReCoVERR Implementation Details
LLaVA-1.5 Vanilla SelPred 8.2 21.5 25.1 33.0 16.5 40.1 59.3 69.9
Vision Tools 9.6 24.4 30.0 38.3 17.3 40.8 61.8 72.1
ReCoVERR 11.3 34.7 45.4 55.9 19.5 44.6 72.7 81.9
@@
4.3.1 ReCoVERR Model-based Tools
ReCoVERR leverages a set of vision tools $\mathbb{M}_{\texttt{Vis}}$ for extracting general visual information. In our experiments, the vision tools $\mathbb{M}_{\texttt{Vis}}$ consists of LVIS Gupta et al. (2019) for object detection, and Qwen-VL Bai et al. (2023) for region captioning. We use FlanT5-XL as a zero-shot sentence paraphraser $\mathcal{M}_{\texttt{QA}\rightarrow\texttt{S}}$ and entailment model $\mathcal{M}_{\texttt{NLI}}$. Since our VLMs use FlanT5-XL as the LLM backbone, our ReCoVERR instantiation does not require additional models. We use GPT-3.5 with temperature 1.0 for $\mathcal{M}_{\texttt{QGen}}$. See Appendix B for prompts.
4.3.2 ReCoVERR Hyperparameters
4.4 Selective Prediction Baselines
We compare ReCoVERR to a Vanilla Selective Prediction baseline Whitehead et al. (2022a), where the model abstains if the VLM prediction confidence $g(a)=\{1\text{ if }\pi_{\texttt{{VLM}}}(a)\leq\gamma_{@r}\text{ else }0\}$, and to a Vision Tools baseline. For Vision Tools, image caption from the VLM, objects detected by LVIS, and region captions from Qwen-VL are provided as evidence to the NLI model $\mathcal{M}_{\texttt{NLI}}$ directly, with no rounds of evidence collection, equivalent to lines 1-5 of Algorithm 1 followed by lines 14-15.
4.5 Evaluation Metrics
We measure risk ($\mathcal{R}$) and coverage ($\mathcal{C}$) (Eqs 1, 2) across system predictions and abstentions. Additionally, we calculate Effective Reliability $\Phi_{c}$ Whitehead et al. (2022a), a metric that rewards correct predictions, assigns a penalty $c$ to incorrect predictions, and assigns zero reward to abstentions. We assign a penalty of $c=1$ for incorrect answers ($\Phi_{1}$). Finally, we define Selective Prediction Recall ($\mathbb{R}_{\texttt{SP}}$), the percentage of correct answers ($\text{Acc}(a)=1$) in the dataset that were answered.
@@ 5 Results and Analysis
At 20% risk tolerance, we find that ReCoVERR improves coverage, effective reliability, and recall over both baselines while staying at the specified risk tolerance. At 10% risk tolerance, ReCoVERR and the Vision Tools baseline tend to slightly overshoot the risk tolerance (by 1-2% for InstructBLIP and LLaVA-1.5, and 3-4% for BLIP2). We further see that ReCoVERR shows large improvements of up to 20% on coverage, 15% on effective reliability and 33% on recall for off-the-shelf BLIP2, calibrated InstructBLIP and LLaVA-1.5.
VLM Method $\mathcal{R}$($\downarrow$) $\Phi_{1}$($\uparrow$) $\mathcal{C}$($\uparrow$) $\mathbb{R}_{\texttt{SP}}$($\uparrow$)
BLIP2 (Off-the-shelf) Vanilla SelPred 07.9 13.5 15.9 21.3
Vision Tools 11.8 20.5 26.6 34.0
ReCoVERR 11.9 21.3 27.8 35.4
BLIP2 (Calibrated) Vanilla SelPred 07.7 13.1 15.5 20.9
Vision Tools 11.3 21.8 27.9 35.9
ReCoVERR 11.8 22.9 29.7 38.1
InstructBLIP (Off-the-shelf) Vanilla SelPred 10.0 31.8 39.5 45.8
Vision Tools 10.8 34.8 43.1 50.2
ReCoVERR 10.7 35.5 45.0 51.9
InstructBLIP (Calibrated) Vanilla SelPred 10.3 32.3 41.1 47.5
Vision Tools 11.2 35.4 44.3 51.1
ReCoVERR 11.0 35.9 45.8 52.6
LLaVA-1.5 Vanilla SelPred 10.4 49.4 62.3 74.5
Vision Tools 11.0 49.7 63.1 75.5
ReCoVERR 11.8 51.5 67.9 79.2
Table 2: Metric results as percentages on the VQAv2 task at a risk tolerance of 10%. We evaluate selective prediction methods on overall system risk ($\mathcal{R}$), effective reliability ($\Phi_{1}$), coverage ($\mathcal{C}$) and recall ($\mathbb{R}_{\texttt{SP}}$). System risks in red exceeded tolerance. Measurements in blue indicate when ReCoVERR outperformed both baselines.
We highlight several takeaways (T*). T1: For the same risk tolerance, vanilla selective prediction has lower coverage and recall for BLIP2 than InstructBLIP and LLaVA-1.5. BLIP2 has not been trained on A-OKVQA, unlike the other two VLMs, and therefore produces more uncertain answers. We further see that ReCoVERR results in larger improvements for BLIP2. T2: Off-the-shelf BLIP2 sees largest risk increase with ReCoVERR. The calibration curves in Figure 3 indicate that off-the-shelf BLIP2 makes overconfident estimates, leading to more unreliable evidences being collected and resulting in a large risk increase for ReCoVERR. T3: At 10% risk, most of the additional risk and recall is caused by Vision Tools, rather than ReCoVERR’s evidence collection. T4: At 20% risk tolerance, ReCoVERR’s evidence collection leads to greatest recall improvement for off-the-shelf BLIP2, calibrated InstructBLIP and LLaVA-1.5.
@@
We also experiment on VQAv2, at 10% risk tolerance. [^4] Since the VLMs have strong performance on VQAv2, we do not perform experiment at 20% risk tolerance. In Table 2, we see that ReCoVERR improves system reliability and recovers more correct answers than the baselines, while crossing the risk threshold by only 1–2%. However, the improvements are smaller compared to the more complex A-OKVQA task, indicating that ReCoVERR may be more useful for complex reasoning tasks. We also observe that, similar to A-OKVQA, the largest improvements are for BLIP2, which wasn’t finetuned on the VQAv2 task.